OpenStudy (anonymous):

Whats the equation of the asymptotes for the hyperbola ((y+4)^2)/(36)-((x-3)^2)/(25) If anyone can understand it?

OpenStudy (anonymous):

This should be your equation \[ \frac{(y+4)^2}{36} -\frac {(x-3)^2}{25}=1 \]

OpenStudy (anonymous):

To find the asymptotes you put replace the minus sing by = sign and you take away the 1. Near Infinity 1 is almost nothing. \[ \frac{(y+4)^2}{36} =\frac {(x-3)^2}{25} \\ \frac{(y+4) }{6} =\pm\frac {(x-3)}{5} \\ y-4 =\pm\frac 6 5{(x-3)}\\ y =4\pm\frac 6 5{(x-3)} \]